# for GitHub Issue ReSolution

### v1

@@ 4.3 RQ 4: How Effective Is Our Planning Process?
The correlation scores are determined by GPT-4 based on a set of criteria defined in Table 3, which spans from a score of $1$, indicating no relevance, to a score of $5$, indicating perfect alignment with high accuracy and completeness. The distribution of these correlation scores is presented in Figure 6, where the vertical axis represents the frequency of each score. Notably, most of the scores are $3$ or above, which implies that the majority of task descriptions are in the right direction concerning planning. Furthermore, the higher scores correlate with a higher probability of issue resolution, indicated by a larger proportion of “resolved” outcomes in scores $4$ and $5$. This signifies that when the generated task description closely aligns with the reference, there is a higher possibility to resolve the issue.
The analysis above demonstrates the effectiveness of both the repository custodian and the project manager agent in the planning process of our framework.
@@ 5 The code changes perfectly align with the task description, fully addressing all specified
4.4 RQ 5: How Effective Is Our Coding Process?
To investigate the effectiveness of the coding process in our framework, we analyze the performance of the developer’s line locating and the issue resolving across instances of varying complexities.
Figure 8 illustrates the distribution of the line location overlap ratio of MAGIS and the baseline models (GPT-4 and Claude-2). The vertical axis quantifies the frequency of occurrences within specific ranges of line location overlap ratios for each group. This visualization reveals that our developer agent frequently attains a line location overlap ratio nearing $1$. Compared with baseline models, the developer agent demonstrates a pronounced preference for higher distribution values close to $1$, and conversely, a reduced preference for lower distribution values near 00. Such a distribution validates the superior performance of MAGIS in line location.
Further analysis is provided in Figure 8, which illustrates the relationship between the line location overlap ratio and the issue resolved ratio within those overlaps. In the figure, the horizontal axis represents the range of overlap ratios for each bar’s corresponding interval, while the height of each bar indicates the resolved ratio for instances within that interval. These resolved ratios correspond to the scale on the left vertical axis. The orange curve represents the cumulative frequency of instances that can be resolved under different overlap ratio thresholds, with the cumulative frequency corresponding to the scale on the right side.
As shown in Figure 8, the right four bars are higher than the five left, which indicates the resolved ratio can increase with the line location overlap. This observation also suggests that accurate line location is helpful for issue resolution. The cumulative frequency curve, shown in orange, provides an additional analysis, indicating the cumulative proportion of issues resolved ratio up to each point along the line location overlap. A steady increase in cumulative frequency accompanies the increase in line location overlap, reinforcing the idea that resolving issues is more successful in areas of high overlap. The slope of the curve’s left half is lower than that of the right half, indicating that the benefits of increasing the overlap ratio are less pronounced at lower overlap ratios than at higher ones. Therefore, the developer agent should prioritize improving its capability of line location.
Moreover, as shown in Table 4, we present a logistic regression analysis that quantifies the correlation between several complexity indices and issue resolution. The results show that GPT-4 has significant negative correlations across the number of files and functions, suggesting that as these indices increase, the likelihood of issue resolution decreases. Conversely, the negative correlations are less pronounced with our model, MAGIS, particularly in the number of files and functions, suggesting mitigation of challenges corresponding to these complexity indices.
Method # Files # Functions # Hunks # Added LoC # Deleted LoC # Changed LoC
GPT-4 $-25.15$* $-25.15$* $-0.06$ $-0.10$ $-0.04$ $-0.21$
MAGIS   $-1.55$*   $-1.55$* $-0.12$* $-0.04$* $-0.06$* $-0.57$*
Figure 7: Comparison of Line Location Overlap between MAGIS (Ours) and Baselines.
Figure 8: Resolved Ratio in Different Line Location Overlap Intervals.
To evaluate the performance of the QA engineer, the ablation experiment is conducted and the results are shown in Table 2. As the table shows, in settings with and without hints, the presence of the QA engineer can increase the resolved ratio by $1.57\%$ and $3.31\%$, respectively. This overall enhancement substantiates the QA engineer’s contribution to improving outcomes.
@@ 4.4 RQ 5: How Effective Is Our Coding Process?
Specifically, there is an issue  [^8] https://github.com/scikit-learn/scikit-learn/issues/9784 from the repository scikit-learn  [^9] https://github.com/scikit-learn/scikit-learn and the reference code change  [^10] https://github.com/scikit-learn/scikit-learn/pull/9288 is shown in Figure 5. During the flow of our framework, the developer firstly modifies the code as shown in Figure 10 but the parameterrandom_state (Line $371$ in the new-version code) of the function kmeans_single is not assigned the right number in seeds. After the erroneous modification was made, the QA engineer identified the mistake and provided feedback. Their commentary highlighted the issue: “This code change modifies the implementation of K-means algorithm and doesn’t seem entirely correct”. They further elaborated, “Running the algorithm just one time could lead to worse results, compared to running it multiple times (n_init times) and choosing the best result, as was originally done” This critique specifically targets the flaw associated with the iterative process (“running times”). With the help of the QA engineer, the developer further revise the code, and the final code change is shown in Figure 10. All of the necessary test cases are passed after applying this code change. Both the ablation study and the case study underscore the QA engineer’s efficacy.
5 Statistics and Discussion
Table 5: Case from scikit-learn (Gold).
Figure 9: Case from scikit-learn (Ours, before review).
Figure 10: Case from scikit-learn (Ours, after review).
5.1 Complexity of Code Changes.
The statistics on the code change for instances with resolved issues are presented in Table 6. Overall, the statistical information of the generated code changes for these instances, such as the average number of code files, functions, hunks, and deleted lines, all differ slightly (not exceeding $0.3$) from the reference solutions written by humans. This indicates that for these instances, the complexity of the code change generated by our framework is similar to that of humans. Furthermore, the maximum values observed in the table reveal that our framework can implement code modifications involving two files, four hunks, and $1,655$ lines modification, with single modifications reaching up to $190$ lines. Results demonstrate the effectiveness of our method in resolving complex issues that need to modify the code file on multiple locations and with long context.
Table 6: The Statistical Analysis of Our Framework on Resolved Instances.
MAGIS Gold
Min Max Avg. Min Max Avg.
# Code Files $1$ $2$ $1.02$ $1$ $2$ $1.04$
# Functions $1$ $2$ $1.02$ $1$ $2$ $1.04$
# Hunks $1$ $4$ $1.45$ $1$ $6$ $1.66$
# Added Lines $1$ $146$ $9.75$ 00 $38$ $4.34$
# Deleted Lines 00 $77$ $5.27$ 00 $115$ $5.16$
Change Start Index $1$ $1,655$ $270.12$ $1$ $1,657$ $256.09$
Change End Index $22$ $1,665$ $301.68$ 00 $1,666$ $315.05$
# Changed Lines $2$ $190$ $15.02$ $1$ $115$ $9.50$
Specifically, the distribution of the number of modified lines for the resolving instances is shown in Figure 12. We observe that the distribution of the number of modified lines in our framework for the solved instances exceeds that of the reference solution, especially in terms of the number of added lines being significantly higher than the reference. Upon manual inspection, we found that the generation results provided by our framework often contained more comment information, which led to an increase in the total number of modified lines. For example, Figure 10 displays the generation result of our framework. Lines $365,368,371,374,383$ in the new version file correspond to the comment for the added code. These natural language descriptions are valuable in actual software evolution [14, 22]. In contrast, Figure 5 shows a human-written solution lacking such explanatory comments, which might disadvantage software maintainers in reading and understanding.
Refer to captionRefer to caption
Figure 11: Distribution of the LoC in the Resolved Instances.
Refer to captionRefer to caption
The statistics on the code change for instances without resolved issues are shown in Table 7. From the table, our framework can generate applicable code changes including up to $13$ files and $28$ hunks, and the location of the modifications can be as far as line $7,150$, with a single modification reaching up to $9,367$ lines. These results suggest that our method has a strong adaptability in generating applicable code changes. However, considering that these code changes have not passed all the potential test cases they could pass, which indicates that there is still room for improvement.
Table 7: The Statistical Analysis of Our Framework on Applied but Not Resolved Instances.
MAGIS Gold
Min Max Avg. Min Max Avg.
# Code Files $1$ $13$ $1.50$ $1$ $18$ $1.61$
# Functions $1$ $13$ $1.50$ $1$ $18$ $1.61$
# Hunks $1$ $28$ $2.52$ $1$ $52$ $3.72$
# Added Lines $1$ $920$ $40.38$ 00 $3,050$ $28.27$
# Deleted Lines 00 $9,160$ $327.27$ 00 $2,975$ $14.51$
Change Start Index $1$ $4,568$ $424.84$ 00 $6,651$ $485.01$
Change End Index $9$ $7,150$ $513.13$ 00 $6,658$ $728.96$
# Changed Lines $1$ $9,367$ $367.65$ $1$ $6,025$ $42.79$
To further analyze the reasons behind the failure of test cases in these instances, we have quantified the distribution of the lengths of code changes in the unresolved instances, as shown in Figure 12. From the figure, we observe that for unresolved instances, the framework tends to delete a larger number of lines while adding fewer lines, in contrast to the distribution of human-written changes. This discrepancy may point to different repair strategies or attitudes towards problem-solving, where the framework presented herein might prefer to reduce errors by removing potentially problematic code, whereas human developers may lean towards adding new code to address issues.
Moreover, a comparison between Table 6 and Table 7 reveals that the latter contains a higher overall number of files, hunks, and changed lines of code. These instances, involving more modification locations, correspond to more complex scenarios. This phenomenon suggests that the performance of our framework in resolving such complex issues requires further enhancement.
Furthermore, the variability in difficulty across different software repositories may influence the effectiveness of code changes. To this end, we compile statistics on the resolved ratios in various software repositories, as shown in Figure 13. From the figure, we observe that there is a significant variation in the resolved ratios across different repositories in our framework. Some repositories have a resolved ratio as high as $40\%$, while others are close to $0\%$. This suggests that the differences among various software such as code structure and coding style can impact the generation and application of the code change.
5.2 Comparison with Devin.
Devin is a novel agent for software development [34], and its performance has also been assessed using the SWE-bench. However, the evaluation dataset employed by Devin differs from the subset used for experiments with GPT-4 reported by the paper of SWE-bench [15]. An analysis of the repository name and pull request ID of each instance reveals that only $140$ instances overlap between the two datasets.
6 Limitation
Prompt.
@@ 6 Limitation
7 Related Work
7.1 Large Language Models
7.2 LLM-Based Multi-Agent System
With the powerful text generation capabilities of LLMs, many researchers [12, 31, 7, 37, 26, 36] have explored the construction of LLM-based Multi-Agent Systems, enabling them to accomplish tasks beyond the capabilities of the LLMs themselves. For example, MetaGPT [12], which simulates the Standardized Operating Procedures (SOPs) of a programming team, completing tasks including definition, design, planning, coding, and testing through constructed roles (e.g., product managers, architects, project managers, etc.). This framework has achieved leading scores on the HumanEval [9] and MBPP [2], outperforming many LLMs, and researchers show its ability to complete a software establishment (e.g., a code repository to play Gomoku game), indicating that a multi-agent framework can better leverage the capabilities of LLMs in code generation tasks. Moreover, Qian et al. [26] designed ChatDev, a virtual development company simulating a human development team, which decomposes requirements into atomic tasks assigned to the developers. Developers mitigate the hallucination that may arise with the LLM through mutual communication and self-reflection mechanisms. Experimental results show that ChatDev can complete the establishment of some small software (averaging no more than $5$ files per software) in a relatively short time (less than $7$ minutes on average). However, these works focus on the transformation from the requirement to software and overlook the code change generation during software evolution which needs not only understanding the requirement but also dealing with the large repository.
8 Conclusion
In conclusion, this paper illuminates the potential of large language models in software evolution, particularly in resolving GitHub issues. Our study identifies the challenges of direct LLM application. To address them, we proposed a novel LLM-based multi-agent framework, MAGIS, enhancing issue resolution through well-designed agents’ collaboration. The superiority of MAGIS on the SWE-bench dataset against popular LLMs highlights its effectiveness, pointing towards a promising direction for integrating LLMs into software evolution workflows. This work not only shows the potential of LLMs in GitHub issue resolution but also explores an LLM-based paradigm for software evolution.

### v2

@@ 4.3 How Effective is Our Planning Process? (RQ 3)
For the Manager agent, we examined the alignment of its generated task descriptions with the reference code change by LLM. Following the study [64], we select GPT-4 as an evaluator to score the correlation between the reference code change and the generated task description. The correlation scores are determined based on a set of criteria defined in Tab. 6. A higher correlation score indicates a better alignment and thus, a more accurate and effective planning direction. The distribution of these correlation scores is presented in Fig. 4. Notably, most of the scores are $3$ or above, implying that the majority of task descriptions are in the right direction concerning planning. Furthermore, the higher scores correlate with a higher probability of issue resolution, indicated by a larger proportion of “resolved” outcomes in scores $4$ and $5$. This signifies that when the generated task description closely aligns with the reference, there is a higher possibility of resolving the issue. The analysis above demonstrates the effectiveness of both the Repository Custodian and the Manager agent in the planning process of our framework.
4.4 How Effective is Our Coding Process? (RQ 4)
To evaluate the effectiveness of the coding process in our framework, we analyze the performance of Developers in locating code lines and resolving issues of different complexity.
Fig. 6 illustrates the distribution of the line locating coverage ratio of MAGIS and the baselines. This visualization reveals that our Developer agent frequently attains a line locating coverage ratio nearing $1$. Compared with baselines, the Developer agent demonstrates a pronounced preference for higher distribution values close to $1$, and conversely, a reduced preference for lower distribution values near 00. Such a distribution validates the superior performance of MAGIS in locating code lines.
Further analysis is provided in Fig. 6 illustrating the relationship between the line locating coverage ratio and the issue resolved ratio within those coverages. As shown in Fig. 6, the right four bars are higher than the five left, which indicates that the resolved ratio can increase with the line locating coverage. This observation also suggests that locating lines accurately is important for issue resolution. The cumulative frequency curve, shown in orange, provides an additional analysis, indicating the cumulative proportion of issues resolved ratio up to each point along the line locating coverage. A steady increase in cumulative frequency accompanies the increase in line locating coverage, reinforcing the idea that resolving issues is more successful in areas of high coverage. The slope of the curve’s left half is lower than that of the right half, indicating that the benefits of increasing the coverage ratio are less pronounced at lower coverage ratios than at higher ones. Therefore, the Developer agent should prioritize improving its capability of locating code lines.
Moreover, as shown in Tab. 3, we present a logistic regression analysis that quantifies the correlation between several complexity indices and issue resolution. The results show that GPT-4 has significant negative correlations across the number of files and functions, suggesting that as these indices increase, the likelihood of issue resolution decreases. Conversely, the negative correlations are less pronounced with our model, MAGIS, particularly in the number of files and functions, suggesting mitigation of challenges corresponding to these complexity indices.
Method # Files # Functions # Hunks # Added LoC # Deleted LoC # Changed LoC
GPT-4 $-25.15$* $-25.15$* $-0.06$ $-0.10$ $-0.04$ $-0.21$
MAGIS   $-1.55$*   $-1.55$* $-0.12$* $-0.04$* $-0.06$* $-0.57$*
Figure 5: Comparison of line locating coverage between MAGIS (Ours) and baselines.
Figure 6: Resolved ratio in different line locating coverage intervals.
To evaluate the performance of the QA Engineer, the ablation experiment is conducted and the results are shown in Tab. 2. As the table shows, in settings with and without hints, the presence of the QA Engineer can increase the resolved ratio by $1.57\%$ and $3.31\%$, respectively. This overall enhancement substantiates the QA Engineer’s contribution to improving outcomes. Furthermore, a case detailed in Appendix I underscores the QA Engineer’s effectiveness.
5 Related Work
Researchers have developed LLM-based multi-agent systems, enabling more complex task completion. For instance, MetaGPT [23, 24] simulates a programming team’s Standardized Operating Procedures (SOPs) and achieves leading scores on benchmarks like HumanEval [12] and MBPP [2]. Similarly, ChatDev [44] functions as a virtual development company, decomposing requirements into atomic tasks and utilizing mutual communication and self-reflection to mitigate LLM hallucinations. While these systems excel in transforming requirements into code, they often overlook the challenges of code change generation during software evolution [25]. GitHub issues include different types of requirements and most of them belong to bug fixing. Previous researchers have proposed methods to localize the bugs [66, 43] and some researchers explored various methods to automatic program repair[58, 7, 56, 3, 60, 54]. The full version of related work can be found in Appendix J.
6 Conclusion
This paper illuminates the potential of LLMs in software development, particularly in resolving GitHub issues. Our empirical study identifies the challenges of direct LLM application. To address the challenges, we propose a novel LLM-based multi-agent framework, MAGIS, enhancing issue resolution through well-designed agents’ collaboration. The superiority of MAGIS on the SWE-bench against popular LLMs highlights its effectiveness, pointing towards a promising direction for integrating LLMs into software evolution workflows.
Appendix A Detailed Explanation in Empirical Study
A.1 Coverage Ratio
The formula for calculating the coverage ratio is Equation 1. As it shows, for each instance of GitHub issue resolution, the range of code change (in terms of the number of lines) in the reference $r$ is represented as a set of intervals ${\bm{L}}_{r}=\{[s_{0},e_{0}],...,[s_{n},e_{n}]\}$, while the line ranges of the generated code change $g$ is ${\bm{L}}_{g}=\{[s_{0}^{\prime},e_{0}^{\prime}],...,[s_{m}^{\prime},e_{m}^{\prime}]\}$, where $s$ and $e$ respectively represent the starting and ending line number of each modification hunk in the file, with $n$ hunks in the reference code change and $m$ hunks in the generated one.
A.2 Observation on Fig. 1
As shown in Fig. 1, we observe that: ① The distribution near the coverage ratio 00 (left side of the figure) is the highest for all three LLMs, indicating that in most cases, the content generated by these models has a very low coverage ratio with the reference in terms of locating code lines. This means that these LLMs are most likely not able to accurately locate code lines that need to be modified in the process of generating the code change. ② In the distribution near the line locating coverage of $1$ (right side of the figure), the three models show a consistent ranking (i.e., Claude-2 $>$ GPT-4 $>$ GPT-3.5) and this ranking is also consistent with the proportion of instances solved by the three models. This phenomenon suggests that the performance of LLMs in generating the code change is probably related to their ability to locate code lines accurately.
A.3 Analysis on Complexity of the Code Change
As shown in Fig. 1, compared with GPT-3.5 and GPT-4, Claude-2 exhibits a different pattern, with much lower negative correlations for the number of files and functions, which indicates that it is a more efficient approach to generate the code change for GitHub issue resolution. However, it also shows significant negative correlations across other indices such as the number of hunks, added lines of code (LoC), deleted LoC, and changed LoC.
Appendix B Kick-off Meeting Example
Figure 7 illustrates a kick-off team meeting. In this meeting, three participants are present: the Manager agent, Oliver CodeLead, and two Developer agents, Django Database Specialist and Alex Rossini. They discuss a specific issue [^1] https://code.djangoproject.com/ticket/30664, assigned tasks, and determine the workflow sequence.
Figure 7: Kick-off meeting to resolve the issue [18].
Appendix C Metrics
The applied ratio indicates the proportion of instances where the code change is successfully generated and can be applied to the existing code repository using Git tools, i.e.,
The recall score versus file number curve is used to measure the effectiveness of locating code files to be modified. The recall score refers to the proportion of files that are successfully located out of all the files that require modification. The formula for calculating the file locating recall score for the $i$-th instance is as follows:
Appendix D Comparison Result on SWE-bench Lite
Recently, some contemporaneous works [39], e.g., AutoCodeRover [61] and SWE-Agent [59], have been proposed for this task. These methods are evaluated using SWE-bench lite, a canonical subset of SWE-bench, which is recommended for evaluation [9]. Considering budget constraints, we conducted experiments on SWE-bench lite to compare with them on the same issues’ resolution.
The experimental results are shown in Tab. 4. MAGIS achieves the highest resolved ratio, $25.33\%$, than other baselines. The performance of MAGIS slightly decreased when evaluated without QA, reaching $23.33\%$, and dropped under the other two ablation settings. This comparative study underscores the robustness of MAGIS, particularly when provided with comprehensive inputs, and highlights the impact of QA and hints on its performance. The results indicate that while new methods like AutoCodeRover and SWE-Agent show promise, MAGIS remains an effective method for GitHub issue resolution.
Table 4: The comparison of overall performance between MAGIS and baselines on SWE-bench lite.
Method AutoCodeRover SWE-Agent MAGIS
Full w/o QA w/o hints w/o (hints, QA)
Resolved $16.11\%$ ($22.33\%$*) $18.00\%$ 25.33% $23.33\%$ $16.67\%$ $16.00\%$
    Note that $16.11$ is the average scores among 3 runs while $22.33$ is under the union of from the 3 runs.
Appendix E Comparison with Devin
Devin is a novel agent for software development [51], and its performance has also been assessed using the SWE-bench. However, the evaluation dataset employed by Devin differs from the subset used for experiments with GPT-4 reported by the paper of SWE-bench [27]. An analysis of the repository name and pull request ID of each instance reveals that only $140$ instances coverage between the two datasets.
Appendix F Statistics on the Generated Code Changes
The statistics on the code change for instances with resolved issues are presented in Tab. 5. Overall, the statistical information of the generated code changes for these instances, such as the average number of code files, functions, hunks, and deleted lines, all differ slightly (not exceeding $0.3$) from the reference solutions written by humans. This indicates that for these instances, the complexity of the code change generated by our framework is similar to that of humans. Furthermore, the maximum values observed in the table reveal that our framework can implement code modifications involving two files, four hunks, and $1,655$ lines modification, with single modifications reaching up to $190$ lines. Results demonstrate the effectiveness of our method in resolving complex issues that need to modify the code file on multiple locations and with long context.
Specifically, the distribution of the number of modified lines for the resolving instances is shown in Fig. 9. We observe that the distribution of the number of modified lines in our framework for the solved instances exceeds that of the reference solution, especially in terms of the number of added lines being significantly higher than the reference. Upon manual inspection, we found that the generation results provided by our framework often contained more comment information, which led to an increase in the total number of modified lines. For example, Fig. 11 displays the generation result of our framework. Lines $365,368,371,374,383$ in the new version file correspond to the comment for the added code. These natural language descriptions are valuable in actual software evolution [26, 35]. In contrast, Fig. 13 shows a human-written solution lacking such explanatory comments, which might disadvantage software maintainers in reading and understanding.
Figure 8: Distribution of the LoC in the resolved instances.
The statistics on the code change for instances without resolved issues are shown in Tab. 5. From the table, our framework can generate applicable code changes including up to $13$ files and $28$ hunks, and the location of the modifications can be as far as line $7,150$, with a single modification reaching up to $9,367$ lines. These results suggest that our method has a strong adaptability in generating applicable code changes. However, considering that these code changes have not passed all the potential test cases they could pass, which indicates that there is still room for improvement.
Figure 10: Case from scikit-learn (ours, after review) for the issue [42].
Figure 11: Case from scikit-learn (ours, before review) for the issue [42].
Figure 12: Case from scikit-learn (gold) [41].
To further analyze the reasons behind the failure of test cases in these instances, we have quantified the distribution of the lengths of code changes in the unresolved instances, as shown in Fig. 9. From the figure, we observe that for unresolved instances, the framework tends to delete a larger number of lines while adding fewer lines, in contrast to the distribution of human-written changes. This discrepancy may point to different repair strategies or attitudes towards problem-solving, where the framework presented herein might prefer to reduce errors by removing potentially problematic code, whereas human developers may lean towards adding new code to address issues.
Moreover, a comparison between the resolved instances and not resolved ones shown in Tab. 5 reveals that the latter contains a higher overall number of files, hunks, and changed lines of code. These instances, involving more modification locations, correspond to more complex scenarios. This phenomenon suggests that the performance of our framework in resolving such complex issues requires further enhancement.
Furthermore, the variability in difficulty across different software repositories may influence the effectiveness of code changes. To this end, we compile statistics on the resolved ratios in various software repositories, as shown in Fig. 13. From the figure, we observe that there is a significant variation in the resolved ratios across different repositories in our framework. Some repositories have a resolved ratio as high as $40\%$, while others are close to $0\%$. This suggests that the differences among various software such as code structure and coding style can impact the generation and application of the code change.
Table 5: The statistical analysis of our framework on resolved and applied but not resolved instances.
Resolved Instances Applied but Not Resolved Instances
MAGIS Gold MAGIS Gold
Min Max Avg. Min Max Avg. Min Max Avg. Min Max Avg.
# Code Files 1 2 1.02 1 2 1.04 1 13 1.50 1 18 1.61
# Functions 1 2 1.02 1 2 1.04 1 13 1.50 1 18 1.61
# Hunks 1 4 1.45 1 6 1.66 1 28 2.52 1 52 3.72
# Added Lines 1 146 9.75 0 38 4.34 1 920 40.38 0 3,050 28.27
# Deleted Lines 0 77 5.27 0 115 5.16 0 9,160 327.27 0 2,975 14.51
Change Start Index 1 1,655 270.12 1 1,657 256.09 1 4,568 424.84 0 6,651 485.01
Change End Index 22 1,665 301.68 0 1,666 315.05 9 7,150 513.13 0 6,658 728.96
# Changed Lines 2 190 15.02 1 115 9.50 1 9,367 367.65 1 6,025 42.79
Appendix G Evaluation on Task Description
Since there is no ground truth for the task descriptions generated by the Manager, we utilize GPT-4 to simulate human evaluation and score each task description based on its corresponding reference code change. Table 6 illustrates the standards used by GPT-4 to assess the correlation between the task description and the code change. The score given by GPT-4 is considered the performance metric for the task description.
@@ 5 The code changes perfectly align with the task description, fully addressing all specified
Appendix H Case Study
Fig. 14 illustrates the detailed process of our framework used to resolve the issue from the Django repository [15] as described in the following ticket  [^4] https://code.djangoproject.com/ticket/30664. To address this issue, two candidate files were identified for modification. Based on the issue description and the candidate files, the Manager defined two file-level tasks. For these tasks, two Developers were assigned: Django Database Specialist (Developer I) and Alex Rossini (Developer II). Following a kick-off meeting attended by both Developers and Managers, the Django Database Specialist commenced work first, followed by Alex Rossini. During the coding phase, Developer I identified the code lines to be modified and generated the new code to replace them. The initial code changes made by Developer I were approved by the QA Engineer. Developer II made three attempts during the coding process. The QA Engineer provided review comments on the first two attempts. Ultimately, both Developers completed their coding tasks, and the merged results from their code changes passed all necessary tests.
Figure 14: Detailed overview of our framework, MAGIS (Kick-off meeting refers to Fig. 7).
Fig. 15 shows a reference issue resolution result, which resolves the issue  [^5] https://code.djangoproject.com/ticket/30255 from the repository Django [15], the human developer modifies four hunks in two files [16]. Despite the presence of modifications in two files, our method focuses on changes in only one file, as shown in Figure 16. Notably, this simpler modification allows the repository to pass all necessary test cases.
Figure 15: Case from Django (gold) [16].
Figure 16: Case from Django (ours) for issue [17].
Appendix I The performance of the QA Engineer Agent
@@
Fig. 13 shows an issue [42] from the repository scikit-learn [40] and the reference code change [41]. During the flow of our framework, the Developer firstly modifies the code as shown in Fig. 11 but the parameterrandom_state (Line $371$ in the new-version code) of the function kmeans_single is not assigned the right number in seeds. After the erroneous modification was made, the QA Engineer identified the mistake and provided feedback. Their commentary highlighted the issue: “This code change modifies the implementation of K-means algorithm and doesn’t seem entirely correct”. They further elaborated, “Running the algorithm just one time could lead to worse results, compared to running it multiple times (n_init times) and choosing the best result, as was originally done”. This critique specifically targets the flaw associated with the iterative process (“running times”). With the help of the QA Engineer, the Developer further revise the code, and the final code change is shown in Fig. 11. All of the necessary test cases are passed after applying this code change.
Appendix J Related Work (Detailed)
J.1 Large Language Models
J.2 LLM-Based Multi-Agent System
With the powerful text generation capabilities of LLMs, many researchers [23, 49, 10, 57, 44, 53, 62] have explored the construction of LLM-based Multi-Agent Systems, enabling them to accomplish tasks beyond the capabilities of the LLMs themselves. For example, MetaGPT [23], which simulates the Standardized Operating Procedures (SOPs) of a programming team, completing tasks including definition, design, planning, coding, and testing through constructed roles (e.g., product managers, architects, project managers, etc.). This framework has achieved leading scores on the HumanEval [12] and MBPP [2], outperforming many LLMs, and researchers show its ability to complete a software establishment (e.g., a code repository to play Gomoku game), indicating that a multi-agent framework can better leverage the capabilities of LLMs in code generation tasks. Moreover, Qian et al. [44] designed ChatDev, a virtual development company simulating a human development team, which decomposes requirements into atomic tasks assigned to the developer agents. Developers mitigate the hallucination that may arise with the LLM through mutual communication and self-reflection mechanisms. Experimental results show that ChatDev can complete the establishment of some small projects (averaging no more than $5$ files per project) in a relatively short time (less than $7$ minutes on average). However, these works focus on the transformation from the requirements to code and overlook the code change generation during software evolution, which requires not only understanding the requirement but also dealing with the large repository.
J.3 Automatic Bug Fixing
GitHub issue resolution is a fundamental aspect of software evolution, with bug fixing being one of the most common scenarios. Fixing bugs involves both bug localization and repair. Previous researchers [66, 43] have developed methods to localize bugs before modifying the code. DreamLoc, proposed by Qi et al. [43], effectively models the characteristics of bug reports and source code files. For automatic program repair, Wong et al. [56] explored a retrieval-based method, while Ye and Monperrus [60] proposed ITER, a generation-based method for handling fault localization re-execution. Additionally, some researchers [54, 55] have combined retrieval techniques with generation models. Recently, Xia et al. [58] demonstrated that directly applying popular LLMs significantly outperforms existing APR methods, showcasing their potential for generating diverse and effective patches. Bouzenia et al. [7] introduced RepairAgent, an autonomous LLM-based agent that plans and executes bug fixes by dynamically interacting with various tools.
Appendix K Limitation
Prompt